# CPCL: Cross-Modal Prototypical Contrastive Learning for Weakly Supervised Text-based Person Re-Identification

### v1

@@ Abstract
Weakly supervised text-based person re-identification (TPRe-ID) seeks to retrieve images of a target person using textual descriptions, without relying on identity annotations and is more challenging and practical. The primary challenge is the intra-class differences, encompassing intra-modal feature variations and cross-modal semantic gaps. Prior works have focused on instance-level samples and ignored prototypical features of each person which are intrinsic and invariant. Toward this, we propose a Cross-Modal Prototypical Contrastive Learning (CPCL) method. In practice, the CPCL introduces the CLIP model to weakly supervised TPRe-ID for the first time, mapping visual and textual instances into a shared latent space. Subsequently, the proposed Prototypical Multi-modal Memory (PMM) module captures associations between heterogeneous modalities of image-text pairs belonging to the same person through the Hybrid Cross-modal Matching (HCM) module in a many-to-many mapping fashion. Moreover, the Outlier Pseudo Label Mining (OPLM) module further distinguishes valuable outlier samples from each modality, enhancing the creation of more reliable clusters by mining implicit relationships between image-text pairs. Experimental results demonstrate that our proposed CPCL attains state-of-the-art performance on all three public datasets, with a significant improvement of 11.58%, 8.77% and 5.25% in Rank@1 accuracy on CUHK-PEDES, ICFG-PEDES and RSTPReid datasets, respectively. The code is available at https://github.com/codeGallery24/CPCL.
Figure 1: Illustration of instance-wise contrastive learning and cross-modal prototypical contrastive learning. (a): It attempts to directly draw the image and text pairs closer together. (b): Each visual and textual instance is assigned a prototype corresponding to modality, according to its pseudo label.
@@ 1 Introduction
Text-based Person Re-Identification (TPRe-ID) is designed to retrieve all images of a pedestrian across multiple cameras based on a textual descriptionLi et al. (2017b); Ding et al. (2021); Zhu et al. (2021). A widely adopted framework for this task involves projecting the image and text input to a common latent space through separate encoders and aligning the features of the same identity in the latent space. The identity annotation is crucial for mitigating intra-class variations in the alignment process. On one hand, the annotation helps mitigate intra-modal feature variations. For example, images of a person could be captured by different cameras with varying illumination, view, and background. On the other hand, identity labels narrow down the cross-modal semantic gaps among input images and texts. However, obtaining identity annotations for the dataset is expensive and time-consuming. In a more practical setting, weakly supervised TPRe-ID Zhao et al. (2021) is considered, where the training set only contains text-image pairs without any identity annotations.
The absence of identity labels consequently introduces intra-class variations at both cross-modal and intra-modal levels. To address this challenge, CMMT Zhao et al. (2021) employed a clustering refinement method to refine pseudo labels, and a Text-IoU method was utilized to create soft labels for cross-modal matching. CAIBCWang et al. (2022a) proposed an instance-level matching method to avoid over-reliance on color information to enhance the performance of weakly supervised TPRe-ID. Although these methods have led to great performance advances in weakly supervised TPRe-ID, they suffer from the following issues. Firstly, the instance-wise contrastive loss treats two samples as a negative pair as long as they are from different pseudo labels and leads to an embedding space where all instances are well-separated, and each instance is locally smoothLi et al. (2021). It ignores the prototypical features of each person, which substantially hindered their further improvements. Secondly, they employed pre-trained unimodal models, such as ResNet-50 and Bi-LSTM as image and text encoders, respectively. However, they lack the necessary underlying alignment capabilities due to training on unimodal datasets. Thirdly, the soft labels generated by the Text-IOU were based solely on the percentage of overlapping noun phrases in the text description, which is unreliable and does not encourage the model to learn the overall semantic information of the caption.
@@
To address the aforementioned challenges, inspired by human cognitive processRoss and Makin (1999); Lakoff (2007) and progress in other visual taskSnell et al. (2017); Li et al. (2021), we propose a Cross-Modal Prototypical Contrastive Learning (CPCL) method to guide the mulit-modal leaning. Intuitively, humans have the metacognitive ability to access knowledge to classify a sample as a certain type and then use their knowledge of this type to determine how to deal with itFleming and Dolan (2012). We propose a novel learning scheme where the visual samples of each person are assigned a visual prototype (and similarly, textual samples are assigned a textual prototype) to leverage this cognitive behaviour mechanism. A prototype is defined as a representative embedding in which samples around a single prototype representation for each modal class. The prototype acts as a bridge, overcoming intra-modal feature variations while mitigating the cross-modal semantic gaps, as illustrated in Fig. 1(b).
@@
Furthermore, we cluster image samples and text samples separately to get a better initialization of the prototypes and assign each sample a pseudo label by DBSCAN algorithmEster et al. (1996). Considering that the prototypical features are constantly changing as the model updates, we propose the Prototypical Multi-modal Memory (PMM) module, which dynamically maintains the prototypical features so that the prototypical features can be kept up-to-date. To mitigate intra-class variation, we propose a Hybrid-level Cross-modal Matching (HCM) module. HCM calculates the Prototypical Contrastive Matching (PCM) loss between instance features and prototypical features of different modalities provided by PMM, along with Instance-level Cross-modal Projection Matching (ICPM) loss in each mini-batch with positive and negative samples obtained according to the pseudo labels. During the training process, we discovered a large number of un-clustered samples. We proposed an Outlier Pseudo Label Mining (OPLM) module to identify valuable outlier samples from each modality, enhancing the clustering quality by mining implicit relationships between image-text pairs. To obtain a better-aligned latent space of visual and textual encoders, we leverage CLIPRadford et al. (2021) as the backbone for our framework.
Our main contributions can be summarized as follows:
  • We propose a Cross-Modal Prototypical Contrastive Learning (CPCL) method designed to mitigate the intra-modal feature variations and cross-modal semantic gaps. This learning scheme is inspired by human cognitive behaviour.
@@
  • We propose a Prototypical Multi-modal Memory (PMM) for Hybrid-level Cross-modal Matching (HCM), which helps CLIP establish the associations between heterogeneous modalities of image-text pairs belonging to the same class.
@@
  • We have conducted extensive experiments that demonstrate our proposed CPCL achieves state-of-the-art results on three public benchmark datasets with up to average 8.53% Rank@1 gains. To the best of our knowledge, our work marks the first introduction of the CLIP pre-trained model to weakly supervised TPRe-ID.
@@
Figure 2: Our proposed Cross-modal Prototypical Contrastive Learning assigned each visual and textual sample a prototype corresponding to modality, according to its pseudo label. PMM: Prototypical Multi-modal Memory module dynamically maintains the prototypical features. HCM: Hybrid-level Cross-modal Matching module consists of Prototypical Contrastive Matching (PCM) loss and Instance-level Cross-modal Projection Matching (ICPM) loss. OPLM: Outlier Pseudo Label MIning module identifies valuable un-clustered samples by mining implicit relationships between image-text pairs.
2 Related Work
2.1 Text-based Person Re-Identification
Text-based Person Re-Identification (TPRe-ID) differs from the classic Re-ID problemZheng et al. (2015); Ge et al. (2020); Zhao et al. (2022) as its objective is to identify pedestrian images based on provided text description. TPRe-ID was originally proposed by LiLi et al. (2017b), which utilized a CNN-RNN network to learn cross-modal features at the global level. Early worksLi et al. (2017a, b); Chen et al. (2018) employed VGGSimonyan and Zisserman (2014) and LSTMHochreiter and Schmidhuber (1997) for learning visual-textual representations. Subsequently, these representations were aligned by using a matching loss function. Building on this, ResNetHe et al. (2016) and BERTKenton and Toutanova (2019) were adopted for feature extraction, along with novel cross-modal matching losses to align global image-text features in a joint embedding space. Fine-grained approachesWu et al. (2021); Chen et al. (2022); Wang et al. (2022a) consider local and global matching relations. With increasing attention to CLIPRadford et al. (2021), more and more effortsYan et al. (2022); Jiang and Ye (2023); Bai et al. (2023) have endeavored to leverage CLIP’s powerful knowledge to establish an effective mapping relationship between images and texts. However, the aforementioned methods require identity annotations that are costly and time-consuming to obtain. To address this limitation, CMMTZhao et al. (2021) introduces weakly supervised text-based person re-identification. CMMT proposed MPLR to mine valuable un-clustered samples while employing Text-IOU to guide the model in learning image-text matching and to mine challenging samples.
2.2 Contrastive Learning
Contrastive learningHadsell et al. (2006) is a self-supervised learning method that learns discriminative feature representations by learning the similarities and dissimilarities among samples, leading to remarkable achievements in various domainsYe et al. (2019); He et al. (2020); Chen et al. (2020); Chen and He (2021); Misra and van der Maaten (2020). These methods are based on instance-wise contrastive learning and need to sample a sufficiently large number of negative instances. Though PCLLi et al. (2021) proposed prototypical contrastive learning to encourage the representation to encode the semantic structure of data, it was evaluated on the unimodal dataset and did not explore the complex relation between different modal samples in the weakly supervised TPRe-ID task. CLIPRadford et al. (2021) aligns images and texts closely in a common embedding space through cross-modal contrastive learning loss. Weakly supervised TPRe-ID is a fine-grained task and CLIP does not perform well if it is directly utilized on this task.
@@ 3.1 Overview
The core of the proposed CPCL is to bring instance-level samples closer to their prototypes and then pull two modal prototypes of the same person together. The complete pipeline of our proposed method is illustrated in Fig. 2. Initially, we input all image-text pairs into the image and text encoder respectively, and extract features by CLIPRadford et al. (2021). Secondly, we use the DBSCANEster et al. (1996) algorithm to cluster images and texts separately, and Outlier Pseudo label Mining (OPLM) will identify valuable outlier samples. Thirdly, initialize Prototypical Multi-modal Memory (PMM) and train modal by Hybrid-level Cross-modal Matching (HCM) loss. Last, the outlier samples that OPLM does not mine in step three will be used to further train the model.
@@ 3.2 Prototpcial Multi-modal Memory
Humans can be capable of classifying a sampler as a certain type and then use their knowledge of this type to determine how to deal with itFleming and Dolan (2012). The proposed cross-modal prototypical contrastive learning simulates this cognitive behavior mechanism by bringing instance-level samples closer to their corresponding prototypes. Considering that the prototypical features are constantly updating as the training process, we propose the Prototypical Multi-modal Memory (PMM) module to dynamically maintain the prototypical features and keep them up-to-date.
3.2.1 Memory Initialization
@@
where $c^{v}_{i}$ represents the cluster centroid feature with pseudo label $i$ in the image memory, while $Y^{v}_{i}$ denotes a set of all image samples with the pseudo label $i$. Furthermore, the symbol |||||$\cdot$||| | ⋅ | | indicates the number of elements in a set. The initialization process for text memory follows the same procedure as image memory.
In conclusion, we perform separate clustering for both images and texts, which yields respective corresponding pseudo labels for each modality.
3.2.2 Memory momentum updating
@@
3.3 Hybrid-level Cross-modal Matching
3.3.1 Protopypical Contrastive Matching
@@
where $f^{v}$ denotes the image feature. The parameters $\tau^{v}$ and $N^{t}$ represent the learnable temperature parameter for the image memory and the number of cluster categories for captions, respectively. Symmetrically, the $PCM_{c}$ loss for text-to-image direction, denoted by $\mathcal{L}^{t}_{c}$ is derived by interchanging the roles of $f^{v}$ with $f^{t}$ and $c^{t}$ with $c^{v}$ within the framework of equation Eq. 3.
In summary, the bi-directional $PCM_{c}$ loss is calculated by:
@@
Similarly, $PCM_{s}$ loss $\mathcal{L}^{s}_{pcm}$ is also bi-directional and composed of $\mathcal{L}^{v}_{s}$ and $\mathcal{L}^{t}_{s}$. For example, replacing the $c^{t+}$ in Eq. 3 with $c^{v+}$ creates $\mathcal{L}^{v}_{s}$. In our paper, cross-modal PCM is adopted by default, $\mathcal{L}_{pcm}=\mathcal{L}_{pcm}^{c}$.
3.3.2 Instance-level Cross-modal Project Matching
In addition to matching at the prototypical level, instance-level matching is also necessary. Inspired by the CMPMZhang and Lu (2018), we applied Instance-level Cross-modal Project Matching (ICPM) loss for alignment at the instance level. Specifically, we calculate the probability of cross-modal matching pairs by the following softmax function:
$p_{i,j}=\frac{\exp\left(f^{v}_{i}\cdot f^{t}_{j}/\tau\right)}{\sum_{k=1}^{N}\exp\left(f^{v}_{i}\cdot f^{t}_{k}/\tau\right)},$ (5)
where footnote $i$ and $j$ denote the indexes of features in a mini-batch. $N$ is the number of image-text pairs. The pseudo cross-modal matching probability is $q_{i,j}=y_{i,j}/\sum_{k=1}^{N}y_{i,k}$. $y_{i,j}=1$ means that $(f^{v}_{i},f^{t}_{j})$ is a matched pair from the same pseudo label, while $y_{i,j}=0$ indicates the unmatched pair. Then the ICPM loss is computed by
$\mathcal{L}_{i2t}^{icpm}=KL(\mathbf{p_{i}}\|\mathbf{q_{i}})=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}p_{i,j}\log(\frac{p_{i,j}}{q_{i,j}+\epsilon}),$ (6)
where $\epsilon$ is a small number to avoid numerical problems. Symmetrically, the ICPM loss from text to image $\mathcal{L}_{t2i}^{icpm}$ can be formulated by exchanging $f^{v}$ and $f^{t}$ in Eqs. 5 and 6. The bi-directional ICPM loss is calculated by:
$\mathcal{L}_{icpm}=\mathcal{L}_{i2t}^{icpm}+\mathcal{L}_{t2i}^{icpm}.$ (7)
@@ 3.3.2 Instance-level Cross-modal Project Matching
$\mathcal{L}_{\textrm{overall}}=\mathcal{L}_{pcm}+\mathcal{L}_{icpm}$ (8)
Figure 3: Illustration of Outlier Pseudo Label Mining (OPLM). 1) For each outlier instance a, we group all the paired samples as a set B and eliminate any outlier samples from B. 2) If B is not empty, we identify the nearest sample set to B within the same modality and denote it as C. 3) Next, we search the set of paired instances D of C in the other modality and remove any outlier samples from D. 4) If D is not empty, we assign the class of the samples closest to a as the class of a.

### v2

@@ Abstract
Weakly supervised text-based person retrieval seeks to retrieve images of a target person using textual descriptions, without relying on identity annotations and is more challenging and practical. The primary challenge is the intra-class differences, encompassing intra-modal feature variations and cross-modal semantic gaps. Prior works have focused on instance-level samples and ignored prototypical features of each person which are intrinsic and invariant. Toward this, we propose a Cross-Modal Prototypical Contrastive Learning (CPCL) method. In practice, the CPCL introduces the CLIP model to weakly supervised text-based person retrieval to map visual and textual instances into a shared latent space. Subsequently, the proposed Prototypical Multi-modal Memory (PMM) module captures associations between heterogeneous modalities of image-text pairs belonging to the same person through the Hybrid Cross-modal Matching (HCM) module in a many-to-many mapping fashion. Moreover, the Outlier Pseudo Label Mining (OPLM) module further distinguishes valuable outlier samples from each modality, enhancing the creation of more reliable clusters by mining implicit relationships between image-text pairs. We conduct extensive experiments on popular benchmarks of weakly supervised text-based person retrieval, which validate the effectiveness, generalizability of CPCL.
Figure 1: Illustration of instance-wise contrastive learning and cross-modal prototypical contrastive learning. (a): It treats two samples as a negative pair as long as they are from different pseudo labels, regardless of their semantic similarity, (b): The prototype acts as a bridge, overcoming intra-modal feature variations and mitigating the cross-modal semantic gaps.
@@ 1 Introduction
Text-based person retrieval is designed to retrieve all images of a pedestrian across multiple cameras based on a textual description(Yan et al. 2022; Jiang and Ye 2023; Bai et al. 2023; Cheng et al. 2024). A widely adopted framework for text-based person retrieval involves projecting the image and text input to a common latent space through separate encoders and aligning the features of the same identity in the latent space. The annotation of the identity is crucial to mitigate the variations intra-class in the alignment process. On the one hand, the annotation helps mitigate intra-modal feature variations. For example, images of a person could be captured by different cameras with varying illumination, view, and background. The annotations provide supervisory information to address intra-modal variation effectively. On the other hand, identity labels narrow down the cross-modal semantic gaps between images and texts. For instance, a textual description should be matched not only with its paired image but also with other images of the same person. However, obtaining identity annotations for the dataset is costly and time consuming. In a more practical setting, the weakly supervised text-based person retrieval (Zhao et al. 2021; Wang et al. 2022; Gong et al. 2024b) is considered, where the training set only contains text-image pairs without any identity annotations.
@@
The absence of identity labels consequently introduces intra-class variations at both cross-modal and intra-modal levels. To address the these challenges, we propose a Cross-Modal Prototypical Contrastive Learning (CPCL) method to guide the multi-modal leaning, inspired by the human cognitive process(Ross and Makin 1999; Lakoff 2007) and the progress in other visual tasks(Snell et al. 2017; Li et al. 2021). Intuitively, humans have the metacognitive ability to access knowledge to classify a sample as a certain type and then use their knowledge of this type to determine how to deal with it(Fleming and Dolan 2012). We propose a noval learning scheme where the visual samples of each person are assigned a visual prototype to leverage this cognitive behavior mechanism. Similarly, textual samples are assigned a textual prototype. Each visual and textual instance is assigned a prototype that corresponds to its pseudo label, which helps the model to learn intra-modal discriminative representations while overcoming intra-modal feature variations while mitigating the cross-modal semantic gaps, as illustrated in Fig. 1(b).
@@
From a technical standpoint, the prototype representations are closely aligned with the class centers obtained via clustering algorithms. To obtain good initial prototype representations, we cluster images and texts separately to get a better initialization of the prototypes and assign each sample a pseudo-label using the DBSCAN algorithm(Ester et al. 1996). Considering that the prototypical representations are constantly changing as the model updates, we propose the Prototypical Multi-modal Memory (PMM) module, which dynamically maintains the prototypical representations so that the prototypical representation can be kept up-to-date. To mitigate intra-class variation, we propose a Hybrid-level Cross-modal Matching (HCM) module. It includes the Prototypical Contrastive Matching (PCM) loss between instance representations and prototypical representations of different modalities provided by PMM, and Instance-level Cross-modal Projection Matching (ICPM) loss in each mini-batch with positive and negative samples obtained according to the pseudo labels. During the training process, we discover a large number of un-clustered samples. We propose an Outlier Pseudo Label Mining (OPLM) module to identify valuable outlier samples from each modality, enhancing the clustering quality by mining implicit relationships between image-text pairs. To obtain a better-aligned latent space of visual and textual encoders, we leverage CLIP(Radford et al. 2021) as the backbone for our framework. Our main contributions can be summarized as follows:
  • We propose a Cross-Modal Prototypical Contrastive Learning (CPCL) framework designed to mitigate the intra-modal representation variations and cross-modal semantic gaps. We conduct extensive experiments on popular benchmarks of weakly supervised text-based person retrieval, which validate the effectiveness, the generalizability of CPCL.
@@
  • We propose a Prototypical Multi-modal Memory (PMM) for Hybrid-level Cross-modal Matching (HCM), which helps CLIP establish the associations between heterogeneous modalities of image-text pairs belonging to the same person.
@@
Figure 2: Our proposed Cross-modal Prototypical Contrastive Learning assigned each visual and textual sample a prototype corresponding to modality. PMM: Prototypical Multi-modal Memory module dynamically maintains the prototypical features. HCM: Hybrid-level Cross-modal Matching module consists of Prototypical Contrastive Matching (PCM) loss and Instance-level Cross-modal Projection Matching (ICPM) loss. OPLM: Outlier Pseudo Label MIning module identifies valuable un-clustered samples by mining implicit relationships between image-text pairs.
2 Rekated Work
2.1 Text-based Person Retrieval
Text-based person retrieval differs from the classic image-based person retrieval (also known as person re-identification) problem(Zheng et al. 2015; Ge et al. 2020; Zhao et al. 2022), and its objective is to identify person images based on a textual description. Text-based person retrieval was originally proposed by Li(Li et al. 2017), which utilized a CNN-RNN network to learn cross-modal representation at the global level and provided a benchmark dataset that sparked extensive research in this domain. As the fields of vision and language have evolved, various innovative approaches have emerged, addressing complex challenges in multimodal contexts(Jiang and Ye 2023; Bai et al. 2023; Yan et al. 2022; Cheng et al. 2024). With increasing attention to CLIP(Radford et al. 2021), more and more efforts(Yan et al. 2022; Jiang and Ye 2023; Bai et al. 2023; Deng et al. 2024; Cheng et al. 2024) have endeavored to leverage CLIP’s powerful knowledge to establish an effective mapping relationship between images and texts. Leveraging labeled data, especially the id annotation, these methods achieve significantly improved performance and greatly contributing to the development of the field. These methods belong to the field of fully supervised learning, but in the practical scenario, the ID annotation is costly and time-consuming.
2.2 Weakly Supervised Text-based Person Retrieval
Research on weakly supervised text-based person retrieval tasks is scarce. There are only a few works on weakly supervised text-based person search. CMMT(Zhao et al. 2021) first proposed the task and proposed MPLR to mine valuable un-clustered samples while employing Text-IOU to guide the model in learning image-text matching and to mine challenging samples. Our Outlier Pseudo Label Mining (OPLM) module is similar to it, but the core difference is that OPLM considers the existence of one-to-one and one-to-many relationships between images and text, not just one-to-one relationships in dataset. Based on weakly supervised text-based person retrieval task, some works(Gong et al. 2024b; Du et al. 2024; Gong et al. 2024a) propose a new task, unsupervised incomplete cross-modal retrieval, where the only difference is that not all images (text) have corresponding text (images) in training dataset. For a fair comparison, we only choose the performance of these methods on the complete dataset. ECCA(Gong et al. 2024b) proposed an enhanced prototype-wise granularity alignment module that achieves alignment of global visual and textual embeddings by utilizing cross-modal unified prototypes for both image and text modality as intermediaries. As a comparison, it is similar to our prototypical multi-modal memory module, but there are three different points: (1) Our prototypes are modality-specific rather than unified; that is, images and texts are each associated with their own set of prototypes. This design choice is motivated by the need to mitigate intra-modal feature variations. (2) The pseudo labels we use are obtained by clustering rather than by the Sinkhorn-Knopp algorithm(Cuturi 2013). (3) Finally, our prototypes are updated using a momentum approach rather than a learnable approach. While CCL(Du et al. 2024) and CANC(Gong et al. 2024a) propose intra-modal prototype-based contrastive learning frameworks, our approach, CPCL, extends this paradigm by incorporating both intra-modal and inter-modal prototype-based contrastive matching losses, thereby enabling more comprehensive representation learning across and within modalities.
@@ 3.1 Overview
In this section, we will elaborate on the proposed CPCL framework. First, we introduce the Prototpcial Multi-model Memory (PMM) moduel in Section 3.2, where we will describe the mechanism of its initialization and update. Then, in Section 3.3, we delve into the Hybrid Cross-modal Matching (HCM) module, which encompasses the Protopypical Contrastive Matching (PCM) and Instance Cross-modal Project Matching (ICPM). Finally, we elaborate the proposed Outlier Pseudo Lable Mining (OPLM) in Section 3.4, which has the refined learning stage and the supplementary learning stage. The overall framework of CPCL is illustrated in Fig. 2.
@@ 3.2 Prototpcial Multi-modal Memory
Humans can be capable of classifying a sampler as a certain type and then use their knowledge of this type to determine how to deal with it(Fleming and Dolan 2012). The proposed cross-modal prototypical contrastive learning simulates this cognitive behavior mechanism by bringing instance-level samples closer to their corresponding prototypes. Considering that the prototypical features are constantly updating as the training process, we propose the Prototypical Multi-modal Memory (PMM) module to dynamically maintain the prototypical features and keep them up-to-date.
Memory Initialization
@@
where $c^{v}_{i}$ represents the prototype feature with pseudo label $i$ in the image memory, while $Y^{v}_{i}$ denotes a set of all image samples with the pseudo label $i$. Furthermore, the symbol |||||$\cdot$||| | ⋅ | | indicates the number of elements in a set. The initialization process for text memory follows the same procedure as image memory.
Memory momentum updating
@@
3.3 Hybrid Cross-modal Matching
Protopypical Contrastive Matching
@@ 3.3 Hybrid Cross-modal Matching
where $f^{v}$ denotes the image feature. The parameters $\tau^{v}$ and $N^{t}$ represent the learnable temperature parameter for the image memory and the number of cluster categories for captions, respectively. Symmetrically, the $PCM_{c}$ loss for text-to-image direction, denoted by $\mathcal{L}^{t}_{c}$ is derived by interchanging the roles of $f^{v}$ with $f^{t}$ and $c^{t}$ with $c^{v}$ within the framework of equation Eq. 3:
$\mathcal{L}^{t}_{c}=-\log\frac{\exp\left(f^{t}\cdot c^{v+}/\tau^{t}\right)}{\sum_{i=1}^{N^{v}}\exp\left(f^{t}\cdot c^{v}_{i}/\tau^{t}\right)}.$ (4)
Take Eq. 3 as an example, we utilize the image feature $f^{v}$ to calculate the loss in text memory. Specifically, based on the paired textual feature $f^{t}$, we can identify a positive sample $c^{t+}$ in the text memory, though maybe the cluster id of the image sample is different from the cluster of the paired caption sample. Moreover, we consider samples from different text memory classes as negative samples. This mechanism helps the model to learn the mapping relationship between images and captions globally. Similarly, Eq. 4 facilitates the model in learning the mapping relationship between text and image. In summary, the bi-directional $PCM_{c}$ loss is calculated by:
@@
Similarly, $PCM_{s}$ loss $\mathcal{L}^{s}_{pcm}$ is also bi-directional and composed of $\mathcal{L}^{v}_{s}$ and $\mathcal{L}^{t}_{s}$:
$\mathcal{L}^{v}_{s}=-\log\frac{\exp\left(f^{v}\cdot c^{v+}/\tau^{v}\right)}{\sum_{i=1}^{N^{t}}\exp\left(f^{v}\cdot c^{v}_{i}/\tau^{v}\right)}$ (6)
and
$\mathcal{L}^{t}_{s}=-\log\frac{\exp\left(f^{t}\cdot c^{t+}/\tau^{t}\right)}{\sum_{i=1}^{N^{v}}\exp\left(f^{t}\cdot c^{t}_{i}/\tau^{t}\right)}.$ (7)
Similarly, the $PCM_{s}$ loss is calculated by:
$\mathcal{L}_{pcm}^{s}=\mathcal{L}^{v}_{s}+\mathcal{L}^{t}_{s}$ (8)
In our paper, cross-modal PCM is adopted by default, $\mathcal{L}_{pcm}=\mathcal{L}_{pcm}^{c}$. It is essential to emphasize that the integration of PCM and PMM is imperative for their collaborative functionality. These two components must be combined to work better.
Instance Cross-modal Project Matching
In addition to matching at the prototypical level, instance-level matching is also necessary. Inspired by the CMPM(Zhang and Lu 2018), we applied Instance Cross-modal Project Matching (ICPM) loss for alignment at the instance level. Specifically, we calculate the probability of cross-modal matching pairs by the following softmax function:
$p^{v2t}_{i,j}=\frac{\exp\left(f^{v}_{i}\cdot f^{t}_{j}/\tau\right)}{\sum_{k=1}^{N}\exp\left(f^{v}_{i}\cdot f^{t}_{k}/\tau\right)},$ (9)
where footnote $i$ and $j$ denote the indexes of features in a mini-batch. $N$ is the number of image-text pairs. The pseudo cross-modal matching probability is $q^{v}_{i,j}=y_{i,j}/\sum_{k=1}^{N}y_{i,k}$. $y_{i,j}=1$ means that $(f^{v}_{i},f^{t}_{j})$ is a matched pair from the same pseudo label, while $y_{i,j}=0$ indicates the unmatched pair. Then the ICPM loss is computed by
$\mathcal{L}_{v2t}^{icpm}=KL(\mathbf{p^{v}_{i}}\|\mathbf{q^{v}_{i}})=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}p^{v}_{i,j}\log(\frac{p^{v}_{i,j}}{q^{v}_{i,j}+\epsilon}),$ (10)
where $\epsilon$ is a small number to avoid numerical problems. Symmetrically, the ICPM loss from text to image $\mathcal{L}_{t2i}^{icpm}$ can be formulated by exchanging $f^{v}$ and $f^{t}$ in Eqs. 9 and 10:
$p^{t2v}_{i,j}=\frac{\exp\left(f^{t}_{i}\cdot f^{v}_{j}/\tau\right)}{\sum_{k=1}^{N}\exp\left(f^{t}_{i}\cdot f^{v}_{k}/\tau\right)},$ (11)
and
$\mathcal{L}_{t2v}^{icpm}=KL(\mathbf{p^{t}_{i}}\|\mathbf{q^{t}_{i}})=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{N}p^{t}_{i,j}\log(\frac{p^{t}_{i,j}}{q^{t}_{i,j}+\epsilon}).$ (12)
The bi-directional ICPM loss is calculated by:
$\mathcal{L}_{icpm}=\mathcal{L}_{v2t}^{icpm}+\mathcal{L}_{t2v}^{icpm}.$ (13)
@@
$\mathcal{L}_{overall}=\mathcal{L}_{pcm}+\mathcal{L}_{icpm}$ (14)
Figure 3: Illustration of Outlier Pseudo Label Mining (OPLM). Different colors indicate different cluster IDs. ➀: For an outlier instance, group all the paired and clustered samples as a set B{b, b’}. ➁: Identify the nearest same modality samples set to B and denote it as C{c, c’}. ➂: Search the set of paired instances D {d, d’} of C in the other modality and remove any outlier samples from D {d, d’}. ➃: Assign the closest sample’s cluster ID to the outlier instance.